# PIE: Parkour with Implicit-Explicit Learning Framework for Legged Robots

### v1

@@ II-A Vision-Guided Locomotion
Figure 2: Overview of the proposed PIE framework. The blue box represents the estimator, which utilizes implicit-explicit estimation to provide the policy network with the estimated vectors. The framework is concurrently optimized with PPO for the actor and critic network and regression for the estimator.
@@ III-B Estimator
The encoder module has two inputs: temporal depth images $\mathbf{d}_{t}^{H_{1}}$ and temporal proprioceptive observations $\mathbf{o}_{t}^{H_{2}}$. The depth observations in the near past are stacked in the channel dimension to form $\mathbf{d}_{t}^{H_{1}}$, while the proprioceptive observations in the near past are concatenated to form $\mathbf{o}_{t}^{H_{2}}$. In this work, we set $H_{1}=2$ and $H_{2}=10$. $\mathbf{d}_{t}^{H_{1}}$ is processed by a CNN encoder to extract depth feature maps, while $\mathbf{o}_{t}^{H_{2}}$ is processed by an MLP encoder to extract proprioceptive features. To facilitate cross-modal reasoning between visual tokens from the 2D depth feature map and proprioceptive features, we employ a shared transformer encoder to further integrate the two modalities. Since the robot is equipped with only one egocentric depth camera, it cannot directly perceive terrain beneath or behind its body visually. Therefore, the outputs of the transformer, which are the depth and proprioceptive features after cross-modal attention processing, are concatenated and fed into a GRU to generate memories of the state and terrain. The outputs of the GRU are then served as the policy network’s input vectors. One of these vectors is a encoded height map estimation vector $\mathbf{z}_{t}^{m}$, which is decoded into $\hat{\mathbf{m}}_{t}$ to reconstruct the high-dimensional height map $\mathbf{m}_{t}$, serving as the explicit estimation of the surrounding terrain. Another vector is a purely latent vector $\mathbf{z}_{t}$. Instead of using a pure auto-encoder, following prior work [5], we adopt a VAE structure to extract features for $\mathbf{z}_{t}$. Namely, we employ the KL divergence as the latent loss. When decoded alongside other vectors with explicit meanings, it reconstructs $\hat{\mathbf{o}}_{t+1}$ to represent the robot’s successor state $\mathbf{o}_{t+1}$, encapsulating implicit information about the robot’s state and surroundings. We use mean-squared-error (MSE) loss between the decoded reconstruction and the ground truth for both explicit and implicit estimations. This constitutes the first level of implicit-explicit estimation.
@@ IV-A Experimental Setup
We designed four sets of ablations to compare against ours, aiming to verify the effectiveness of PIE’s dual-level implicit-explicit estimation:

### v3

@@ II-A Vision-Guided Locomotion
Figure 2: Overview of the proposed PIE framework. The grey box represents the estimator, which utilizes implicit-explicit estimation to provide the policy network with the estimated vectors. The framework is concurrently optimized with PPO for the actor and critic network and regression for the estimator.
@@ III-B Estimator
The encoder module has two inputs: temporal depth images $\mathbf{d}_{t}^{H_{2}}$ and temporal proprioceptive observations $\mathbf{o}_{t}^{H_{1}}$. The depth observations in the near past are stacked in the channel dimension to form $\mathbf{d}_{t}^{H_{2}}$, while the proprioceptive observations in the near past are concatenated to form $\mathbf{o}_{t}^{H_{1}}$. In this work, we set $H_{1}=10$ and $H_{2}=2$. $\mathbf{d}_{t}^{H_{2}}$ is processed by a CNN encoder to extract depth feature maps, while $\mathbf{o}_{t}^{H_{1}}$ is processed by an MLP encoder to extract proprioceptive features. To facilitate cross-modal reasoning between visual tokens from the 2D depth feature map and proprioceptive features, we employ a shared transformer encoder to further integrate the two modalities. Since the robot is equipped with only one egocentric depth camera, it cannot directly perceive terrain beneath or behind its body visually. Therefore, the outputs of the transformer, which are the depth and proprioceptive features after cross-modal attention processing, are concatenated and fed into a GRU to generate memories of the state and terrain. The outputs of the GRU are then served as the policy network’s input vectors. One of these vectors is a encoded height map estimation vector $\mathbf{z}_{t}^{m}$, which is decoded into $\hat{\mathbf{m}}_{t}$ to reconstruct the high-dimensional height map $\mathbf{m}_{t}$, serving as the explicit estimation of the surrounding terrain. Another vector is a purely latent vector $\mathbf{z}_{t}$. Instead of using a pure auto-encoder, following prior work [5], we adopt a VAE structure to extract features for $\mathbf{z}_{t}$. Namely, we employ the KL divergence as the latent loss. When decoded alongside other vectors with explicit meanings, it reconstructs $\hat{\mathbf{o}}_{t+1}$ to represent the robot’s successor state $\mathbf{o}_{t+1}$, encapsulating implicit information about the robot’s state and surroundings. We use mean-squared-error (MSE) loss between the decoded reconstruction and the ground truth for both explicit and implicit estimations. This constitutes the first level of implicit-explicit estimation.
@@ IV-A Experimental Setup
We designed five sets of ablations to compare against ours, aiming to verify the effectiveness of PIE’s dual-level implicit-explicit estimation: